OpenStudy (anonymous):

Find the excluded value of x for the polynomial expression: . x-8/8x+1 8 -8 1/8 -1/8

OpenStudy (amistre64):

why would we need to exclude any value?

OpenStudy (anonymous):

idk

OpenStudy (amistre64):

ok, tell me then; what is 5/0 ?

OpenStudy (amistre64):

if we try to solve that we get: 5/0 = n 5 = 0*n ; how could we possibly work that out?

OpenStudy (anonymous):

i dont kno thats y i need help

OpenStudy (amistre64):

this is the help

OpenStudy (amistre64):

dividing by zero is a mathical nono. so, we need to exclude anything that would make the denominator equal to zero

OpenStudy (amistre64):

when does: 8x-1 = 0 ?

OpenStudy (amistre64):

err, 8x + 1 = 0 that is

OpenStudy (anonymous):

when its a negative

OpenStudy (amistre64):

good start, x would have to be a negative number at least

OpenStudy (amistre64):

you have two negatives to choose from; if anything you can plug them in to see what fits

OpenStudy (anonymous):

so it would be -1/8

OpenStudy (anonymous):

so it would be -1/8

OpenStudy (amistre64):

-1/8 is correct :)